OpenStudy (anonymous):

solve for x. x/3+1/2=5/6 a. 3 b. -6 c. 6 d 1

OpenStudy (anonymous):

x/3=2/6 x=1 so d

OpenStudy (anonymous):

x/3+1/2=5/6 x/3=2/6 x=1

OpenStudy (anonymous):

d